OpenStudy (anonymous):

What is the solution of the system? 3x+3y=10 -9x-9y=-30

OpenStudy (anonymous):

I multiplied the first eq. by 3 so I could eliminate a variable. 3(3x+3y=10) 9x+9y=30 A tried adding eq. 1 and 2 to eliminate a variable but then it would just be 30

OpenStudy (nirmalnema):

u will get 18y = 60

OpenStudy (anonymous):

how did you get that though?

OpenStudy (nirmalnema):

9x+9y=30 -9x+9y=30 adding them 18y=60 -9xand 9x will get cancle y=10/3

OpenStudy (anonymous):

But the second equation is : -9x-9x=-30

OpenStudy (anonymous):

you wrote it as: -9x+9y=30

OpenStudy (anonymous):

And sorry, I meant -9x-9y=-30

OpenStudy (nirmalnema):

ohh sorry i havent seen.....i only seen your multiplication

OpenStudy (anonymous):

I don't know what I am doing wrong. b/c after looking at the equation again, all the numbers would cancel out

OpenStudy (nirmalnema):

then there is no solution of these two lines because these are the eq of two parallel lines

OpenStudy (anonymous):

I see, I don't like these problems haha they confuse me. I wish they all had an answer

hero (hero):

Hmmm now look what happens if I just divide both sides of -9x-9y=-30 by -3 -3(3x + 3y) = -3(10) 3x + 3y = 10 Wow, I get the first equation. Must mean that both equations are the same. Thus infinite solutions.

OpenStudy (nirmalnema):

ohh yup i m sorry these are coincident lines

OpenStudy (anonymous):

@Hero can you help me understand what you did? and how you got that answer?

hero (hero):

Before using substitution or elimination method, I always look to see if I can reduce the system first.

hero (hero):

If one of your equations in a system look like 4x + 8y = 12 Clearly that can be reduced to x + 2y = 3 Since each coefficient is divisible by 4

hero (hero):

A better way to understand that is this You have 4x + 8y = 12 4x + 8y factors to 4(x + 2y) 12 factors to 4*3 So 4x + 8y = 12 can be re-written as 4(x + 2y) = 4*3 Now you divide both sides by 4 to get x + 2y = 3

hero (hero):

You should be able to understand one of the two methods of reducing equations.

hero (hero):

When they say "simplify first" before solving, that applies to systems of equations as well.

OpenStudy (anonymous):

alright, they didn't say that in my lessons or in my tutor sessions. That is helpful.

hero (hero):

There's all kinds of little tricks and methods you can use. To be honest, the Indians know all of the cyptic methods of approaching math that isn't taught here. If you want to really learn some solving methods get your hands on some of those Indian books like Quantitative Aptitude and learn the methods in it.

hero (hero):

Of course, by "here" I assume you are American. I could be wrong though.

OpenStudy (anonymous):

I am indeed an American haha thank you so much for helping. and thank you nirmalnema for helping also.

hero (hero):

Simplifying first is sort of an unspoken rule of algebra. It is an implied step. Meaning, no one is going to explicitly tell you. It is just something you have to "see" for yourself during the solving process.

hero (hero):

If you notice that something can be reduced on both sides of the equation or by cancelling factors of one, or by cancelling roots and exponents, then go ahead and do it. It will only make your life easier.

OpenStudy (anonymous):

alright, thank you for telling me that haha it does make things much easier